laplace transform
1. Find the laplace transform of the give functiong! so tho s e fit) and ft). I I sinut cost too I tetsinst tio
Answer #1

Teinwlcornt to I by Defa of Luhluce franalugmation Sinul coral) at E SE (2 8imal corwa) di Ett Rimatul) di * L Et sinaut) t4(4641) = sem er einst da 28+) +8inst dt B = F(8+1) (by first shifting the semillett(+1) = f(S-a)) Now, 7 - Fi 8) = 1ltsinst)